# Quasi Zigzag Persistence: A Topological Framework for Analyzing Time-Varying Data

### v2

@@ 3.1 Stability of Zz-Gril
We prove the stability of Zz-Gril by showing that its perturbation is bounded by the interleaving distance between two quasi zigzag persistence modules. The definition of interleaving distance (see Definition A.3 in Appendix A) between two zigzag persistence modules can be extended to quasi zigzag persistence modules.
@@
Let $(\mathbb{ZZ}\times\mathbb{Z})$ denote the collection of all subposets in $\mathbb{ZZ}\times\mathbb{Z}$ such that their corresponding subposets in $\mathbb{Z}^{2}$ are intervals.
@@
$d_{\mathcal{E}}^{\mathcal{L}}(M,N)\coloneqq\inf\left\{\epsilon\geq 0:\forall\boxed{\mathbf{p}}^{2}_{\delta}\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\text{ and }\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\right\}.$
@@
Let $M$ and $N$ be two quasi zigzag persistence modules. Given appropriate $\mathbf{p},k$, let the Zz-Gril of $M$ and $N$ over $\boxed{\mathbf{p}}^{2}_{\delta}$ be $\lambda^{M}$ and $\lambda^{N}$ respectively. Then,
$|\lambda^{M}(\mathbf{p},k)-\lambda^{N}(\mathbf{p},k)|=d_{\mathcal{E}}^{\mathcal{L}}(M,N)\leq d_{\mathcal{I}}(M,N).$
Corollary 3.6 (Stability).
Let $M$ and $N$ be two quasi zigzag persistence modules. Then,
$||\lambda^{M}-\lambda^{N}||_{\infty}\leq d_{\mathcal{I}}(M,N).$
We refer the reader to Appendix A for the proofs of Proposition 3.4 and Theorem 3.5.
@@ 4 Algorithm
We discuss the details of computing Zz-Gril in this section. We begin by proving that generalized rank over a subposet in $\mathbb{ZZ}\times\mathbb{Z}$ can also be computed by computing the number of full bars in the zigzag filtration along the “boundary” of the subposet. To be precise, we introduce the following concepts now. Let $P$ be any finite connected poset. We say a point $q\in P$ covers a point $p\in P$, denoted $p\prec q$ or $q\succ p$, if $p\leq q$ and there is no $r\in P$, $r\not=p,q$, so that $p\leq r\leq q$. A covering path $p\sim q$ between $p$ and $q$ is either a set of points $\{p=p_{0}\prec p_{1}\prec\cdots\prec p_{\ell}=q\}$ or $\{p=p_{0}\succ p_{1}\succ\cdots\succ p_{\ell}=q\}$. We define some special points in $P$. The set of minima in $P$ are the points $P_{min}=\{p\in P\ \,|\,\not\exists r\not=p\in P\mbox{ where }r\leq p\}$. Similarly, define the set of maxima $P_{max}$. For any two points $p,q\in P$, let $p\vee q$ and $p\wedge q$ denote their least upper bound (lub) and largest lower bound (llb) in $P$ if they exist.
Now consider a subposet $I\subseteq\mathbb{ZZ}\times\mathbb{Z}$ whose counterpart $I^{\mathbb{Z}^{2}}$ with $\mathbb{Z}^{2}$ ordering becomes an interval. Since $I^{\mathbb{Z}^{2}}\subset\mathbb{Z}^{2}$, its points can be sorted with increasing $x$-coordinates in $\mathbb{Z}^{2}$. Let $I^{\mathbb{Z}^{2}}_{min}=\{p_{0},\ldots,p_{s}\}$ and $I^{\mathbb{Z}^{2}}_{max}=\{q_{0},\ldots,q_{t}\}$ be the set of minima and maxima respectively sorted according to their $x$-coordinates. The lower fence and upper fence of $I^{\mathbb{Z}^{2}}$ are defined as the paths
@@
Next, we appeal to certain results in category theory to claim a result analogous to Dey et al. (2024) that helps computing the generalized ranks with zigzag persistence modules. Treating $I$ as a category with points in $I$ as objects and relations as morphisms, we get $\partial_{L}I$ and $\partial_{U}I$ as subcategories. Then, we get functors $F_{L}:\partial_{L}I\rightarrow I$ induced by inclusion and $F_{U}:I\rightarrow\partial_{U}I$ induced by projection which have some special properties.
@@
We call a functor $F:P^{\prime}\rightarrow P$ for $P^{\prime}\subseteq P$ initial if $P^{\prime}$ is initial. Similarly, we call $F:P\rightarrow P^{\prime}$ terminal if $P^{\prime}$ is terminal. The following result connecting initial (terminal) functor to the limit (colimit) is well known, [Chapter 8] Riehl (2014).
Proposition 4.2.
Let $M:P\rightarrow{\bf vec}$ be a persistence module and $F_{L}:P^{\prime}\rightarrow P$ and $F_{U}:P\rightarrow P^{\prime\prime}$ be initial and terminal functors. Then, there is an isomorphism $\phi:\mathsf{lim}{M}\rightarrow\mathsf{lim}{M|_{P^{\prime}}}$ from the limit of $M$ to the limit of the restricted module $M|_{P^{\prime}}$. Similarly, we have an isomorphism for colimits, $\psi:\mathsf{colim}{M|_{P^{\prime\prime}}}\rightarrow\mathsf{colim}{M}$.
We observe that the functors $F_{L}:\partial_{L}I\rightarrow I$ and $F_{U}:I\rightarrow\partial_{U}I$ mentioned earlier are initial and terminal respectively allowing us to use Proposition 4.2.
@@
In the second case, the downset $\downarrow p$ consists of three sets of points $Y^{p^{-}}$, $Y^{p}$, and $Y^{p^{+}}$ that are vertically below $p^{-}$, $p$, and $p^{+}$ respectively. Each of $Y^{p^{-}}$, $Y^{p}$ and $Y^{p^{+}}$ intersects $\partial_{L}I$ in a connected poset. They can be disconnected as a whole only if there is a point $q\in\partial_{L}I$ so that $p^{-}_{x}<q_{x}<p_{x}$ or $p_{x}<q_{x}<p^{+}_{x}$. But, neither is possible because $p^{-}$ and $p^{+}$ are points in $I$ covered by $p$. ∎
Following proposition is immediate.
@@
Now, we prove the following result which extends Theorem 3.1 in Dey et al. (2024) from $2$-parameter persistence modules to the quasi zigzag persistence modules.
@@
Let $M:P\rightarrow{\bf vec}$ be a quasi zigzag persistence module and $I\subseteq P$ be a finite subposet so that $I^{\mathbb{Z}^{2}}$ is an interval. Then, $\mathsf{rk}^{M}(I)=\mathsf{rk}^{M}(\partial I)$.
@@
Figure 3: A worm $I$ (shaded) on a quasi zigzag bi-filtration where the direction of the inclusion on horizontal arrows are shown at the bottom and on the vertical arrows on left. The worm centered at (3,3) has width 1. The part of the boundary colored green is $\partial_{L}I$ connecting the minima shown as green points and the part in purple is $\partial_{U}I$ connecting the maxima shown as purple points. The boundary cap is shown in orange which runs parallel to the boundary for a portion of it. Refer to Figure 5 for the zigzag filtration along the boundary cap of the worm. A sequence of graphs with $T=3$ time steps is shown with the corresponding graph filtrations: $\mathcal{F}_{t_{1}},\mathcal{F}_{t_{2}},\mathcal{F}_{t_{3}}$ each with $L=5$ levels. The filtration of the unions are encircled by ovals. Zigzag filtration at the topmost level $\mathcal{Z}_{L}$ is shown in a rectangular box.
@@ 4.1 Building Quasi Zigzag Bi-filtration
A sequence of $T$ number of graphs is converted into a zigzag filtration $\mathcal{Z}_{L}$ of length $12T-$ 1 by the above procedure because every consecutive pair of graphs introduces a union in between. For each graph in the zigzag filtration thus obtained, we construct its graph filtration based on edge-weights; see e.g. Dey & Wang (2022); Edelsbrunner & Harer (2010). We ensure that the length of each graph filtration remains the same by considering the sublevel sets of exactly $L$ levels. The zigzag filtration at the highest level, $\mathcal{Z}_{L}$, can be pulled back to each lower level $l$. This ensures that we get a zigzag filtration $\mathcal{Z}_{l}$ at each level $1\leq l<L$ of the filtration. This gives us a quasi zigzag bi-filtration. Refer to Figure 3 for an illustation of a quasi zigzag bi-filtration.
@@
Figure 6: Pipeline used for the experiments.
4.2 Computing Zz-Gril
@@ 5 Experiments
In this section, we report the results of testing Zz-Gril on various datasets. We begin by giving a detailed description about the experimental setup. Then, we give a brief description of the datasets, followed by the experimental results. We use the benchmark UEA multivariate time-series Bagnall et al. (2018) datasets to test Zz-Gril on multivariate time series data to show that Zz-Gril can be applied to datasets from various domains. Further, we test Zz-Gril on a targeted application of sleep-stage classification by performing experiments on ISRUC-S3 Khalighi et al. (2016) dataset. In all these experiments, we augment the topological information captured by Zz-Gril to one of the specifically tailored machine learning methods on the respective datasets and compare.
@@ 5.1 Experimental Setup
Given a sample of multivariate time-series data with $m$ time-series, we splice each time-series into a sequence of time-series, each of length $w$. This splicing is done by a moving window of width $w$, where consecutive windows have an overlap of $\lambda$. We calculate the Pearson correlation coefficient Pearson & Lee (1903) between time-series in each window. We construct a graph for each window, where each time-series (of length $w$) is a node, and edges with the Pearson correlation values as weights connect them. Thus, we get a complete graph with $m$ nodes. We select the top $k$ percentile of these edges to build the final graph in each window. This way we obtain a sequence of graphs. The topological information of this sequence of graphs encodes the evolution of correlation between time series. Alternately, we can also track the evolving time-series by converting it into a sequence of point clouds. From the sequence of time-series described above, we consider each time-series of length $w$ as a point in $\mathbb{R}^{w}$. Thus, if we have $m$ time series, we have $m$ points in $\mathbb{R}^{w}$ in each window. Thus, we get a sequence of point clouds in $\mathbb{R}^{w}$. This sequence of point clouds tracks the evolution of each time-series, and a Vietoris-Rips like construction on this point cloud tracks the evolution of the interaction between time-series. Refer to Figure 6 for an illustration. Refer to Appendix LABEL:app:exp for hyperparameters.
@@ 5.2 UEA Multivariate Time Series Classification
UEA Multivariate Time Series Classification (MTSC) Bagnall et al. (2018) archive comprises of real-world multivariate time series data and is a widely recognized benchmark in time series analysis. The UEA MTSC collection encompasses a diverse range of application domains such as healthcare (ECG or EEG data), motion recognition (recorded using wearable sensors). See Table LABEL:tab:app:dataset_info in Appendix LABEL:app:exp.
The datasets are preprocessed and split into training and testing sets. For this set of experiments, we convert the multivariate time series into a sequence of point clouds and compute Zz-Gril. We choose 5 datasets which have at least 7 multivariate time series. This ensures that each point cloud, in the sequence of point clouds, has at least 7 points, giving meaningful topological information. Refer to Figure 7 for a visualization of Zz-Gril on FingerMovements dataset processed as a sequence of point clouds.
In Table 2, we compare the performance of augmenting Zz-Gril to two specifically tailored machine learning models for multivariate time series classification on UEA MTSC datasets, TapNet Zhang et al. (2020) and TodyNet Liu et al. (2024). We can see from the table that Zz-Gril is adding meaningful topological information on most datasets which is seen as an improvement in performance. We picked these two models for augmenting Zz-Gril because of two reasons: (i) availability and ease-of-use of codebase, (ii) these models already have good performance on UEA MTSC Datasets and we wanted to test the additional value of the topological information that Zz-Gril adds. In order to do a comprehensive evaluation of Zz-Gril framework, we compare Zz-Gril in isolation with other methods Li et al. (2021); Tang et al. (2022); Karim et al. (2019); Schäfer & Leser (2017) on UEA MTSC task. We report the results in Table 1. We see that Zz-Gril has comparable performance. Here, we would like to remind the reader that Zz-Gril is a general framework which does not need to be tailored for these datasets in particular. The aim for these experiments is to show that topological information is meaningful and is captured by Zz-Gril. Further, to show that we need both the spatial and the temporal information to be captured together, we compare Zz-Gril with standard zigzag persistence at various spatial scales. We report the results in Table 3. We can see that Zz-Gril performs better than standard zigzag persistence even across different scales.
@@
We perform another set of experiments where we compare the performance of converting the multivariate time series into sequence of point clouds versus converting it into sequence of graphs for extracting topological information. We report the results in Table 5. We can see that both of these approaches are viable choices.
We report the computation times in Table 4 in. We can see from the table that Zz-Gril is, indeed, practical to use.
@@
Dataset TodyNet + Zz-Gril point clouds TodyNet + Zz-Gril graphs
FingerMovements 0.660 0.680
NATOPS 0.961 0.945
SelfRegulationSCP2 0.600 0.594

### v3

@@ 3.1 Stability of Zz-Gril
We prove the stability of Zz-Gril by showing that its perturbation is bounded by the interleaving distance between two quasi zigzag persistence modules. The definition of interleaving distance, $d_{\mathcal{I}}(M,N)$, (see Definition A.3 in Appendix A) between two zigzag persistence modules $M,N$ can be extended to quasi zigzag persistence modules.
@@
Let I(×)($\mathbb{ZZ}\times\mathbb{Z}$) denote the collection of all subposets in $\mathbb{ZZ}\times\mathbb{Z}$ such that their corresponding subposets in $\mathbb{Z}^{2}$ are intervals. For $\epsilon\in\mathbb{Z}_{+}$, the $\epsilon$-thickening of $I$ is defined as
[…]
@@
$\begin{split}d_{\mathcal{E}}^{\mathcal{L}}(M,N)\coloneqq\inf\limits_{\epsilon\geq 0}\{&\forall\boxed{\mathbf{p}}^{2}_{\delta}\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\\ &\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\text{ and }\\ &\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\}.\end{split}$
Note that $\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}$ contains the $\epsilon$-thickening of $\boxed{\mathbf{p}}^{2}_{\delta}$.
@@
Let $M$ and $N$ be two quasi zigzag persistence modules. Let $\lambda^{M}$ and $\lambda^{N}$ denote the Zz-Gril functions of $M$ and $N$ respectively. Then,
$\lambda^{M}-\lambda^{N}||_{\infty}=d_{\mathcal{E}}^{\mathcal{L}}(M,N)\leq d_{\mathcal{I}}(M,N).$
We refer the reader to Appendix A for all the proofs.
@@ 4 Algorithm
We discuss the details of computing Zz-Gril in this section. We begin by proving that generalized rank over a finite subposet in $\mathbb{ZZ}\times\mathbb{Z}$ can also be computed by computing the number of full bars in the zigzag filtration along the “boundary” of the subposet. To be precise, we introduce the following concepts now. Let $P$ be any finite connected poset. We say a point $q\in P$ covers a point $p\in P$, denoted $p\prec q$ or $q\succ p$, if $p\leq q$ and there is no $r\in P$, $r\not=p,q$, so that $p\leq r\leq q$. A covering path $p\sim q$ between $p$ and $q$ is either a set of points $\{p=p_{0}\prec p_{1}\prec\cdots\prec p_{\ell}=q\}$ or $\{p=p_{0}\succ p_{1}\succ\cdots\succ p_{\ell}=q\}$. We define some special points in $P$. The set of minima in $P$ are the points $P_{min}=\{p\in P\ \,|\,\not\exists r\not=p\in P\mbox{ where }r\leq p\}$. Similarly, define the set of maxima $P_{max}$. For any two points $p,q\in P$, let $p\vee q$ and $p\wedge q$ denote their least upper bound (lub) and largest lower bound (llb) in $P$ if they exist.
Now consider a finite subposet $I\subseteq\mathbb{ZZ}\times\mathbb{Z}$ whose counterpart $I^{\mathbb{Z}^{2}}$ with $\mathbb{Z}^{2}$ ordering becomes an interval. Since $I^{\mathbb{Z}^{2}}\subset\mathbb{Z}^{2}$, its points can be sorted with increasing $x$-coordinates in $\mathbb{Z}^{2}$. Let $I^{\mathbb{Z}^{2}}_{min}=\{p_{0},\ldots,p_{s}\}$ and $I^{\mathbb{Z}^{2}}_{max}=\{q_{0},\ldots,q_{t}\}$ be the set of minima and maxima respectively sorted according to their $x$-coordinates. The lower fence and upper fence of $I^{\mathbb{Z}^{2}}$ are defined as the paths
@@
Next, we appeal to certain results in category theory to claim a result analogous to Dey et al. (2024) that helps computing the generalized ranks with zigzag persistence modules.
@@
Definition 4.2.
We call a functor $F:P^{\prime}\rightarrow P$ for $P^{\prime}\subseteq P$ initial if $P^{\prime}$ is initial. Similarly, we call $F:P\rightarrow P^{\prime}$ terminal if $P^{\prime}$ is terminal.
Treating $I$ as a category with points in $I$ as objects and relations as morphisms, we get $\partial_{L}U$ and $\partial_{U}I$ as subcategories. Then, we get functors $F_{L}\colon\partial_{L}I\to I$ induced by inclusion and $F_{U}\colon I\to\partial_{U}I$ induced by projection.
@@
In the second case, the downset $\downarrow p$ consists of three sets of points $Y^{p^{-}}$, $Y^{p}$, and $Y^{p^{+}}$ that are vertically below $^{-}, pp$, and $p^{+}$ respectively. Each of $Y^{p^{-}}$, $Y^{p}$ and $Y^{p^{+}}$ intersects $\partial_{L}I$ in a connected poset. They can be disconnected as a whole only if there is a point $q\in\partial_{L}I$ so that $p^{-}_{x}<q_{x}<p_{x}$ or $p_{x}<q_{x}<p^{+}_{x}$. But, neither is possible because $p^{-}$ and $p^{+}$ are points in $I$ covered by $p$. ∎
The following result connecting initial (terminal) functor to the limit (colimit) is well known, [Chapter 8] Riehl (2014).
Proposition 4.4.
Let $M:I\rightarrow{\bf vec}$ be a persistence module and $F_{L}:\partial_{L}I\rightarrow I$ and $F_{U}:I\rightarrow\partial_{U}I$ be initial and terminal functors. Then, there is an isomorphism $\phi:\mathsf{lim}{M}\rightarrow\mathsf{lim}{M|_{\partial_{L}I}}$ from the limit of $M$ to the limit of the restricted module $M|_{\partial_{L}I}$. Similarly, we have an isomorphism for colimits, $\psi:\mathsf{colim}{M|_{\partial_{U}I}}\rightarrow\mathsf{colim}{M}$.
@@
Proof.
The proof is immediate from the definitions. ∎
For computations, we bank on the following result which extends Theorem 3.1 in Dey et al. (2024) from $2$-parameter persistence modules to the quasi zigzag persistence modules.
@@
Let $M$ be a quasi zigzag persistence module and $I$ be a finite interval in the corresponding quasi zigzag poset. Then, $\mathsf{rk}^{M}(I)=\mathsf{rk}^{M}(\partial I)$.
@@ 4.1 Building Quasi Zigzag Bi-filtration
A sequence of $T$ number of graphs is converted into a zigzag filtration $\mathcal{Z}_{L}$ of length 2T12T-1 by the above procedure because every consecutive pair of graphs introduces a union in between. For each graph in the zigzag filtration thus obtained, we construct its graph filtration based on edge-weights; see e.g. Dey & Wang (2022); Edelsbrunner & Harer (2010). We ensure that the length of each graph filtration remains the same by considering the sublevel sets of exactly $L$ levels. The zigzag filtration at the highest level, $\mathcal{Z}_{L}$, can be pulled back to each lower level $l$. This ensures that we get a zigzag filtration $\mathcal{Z}_{l}$ at each level $1\leq l<L$ of the filtration. This gives us a quasi zigzag bi-filtration. Refer to Figure 4 for an illustation of a quasi zigzag bi-filtration.
@@
4.2 Computing Zz-Gril
Figure 4: A worm II (shaded) on a quasi zigzag bi-filtration where the direction of the inclusion on horizontal arrows are shown at the bottom and on the vertical arrows on left. The worm centered at (3,3) has width 1. The part of the boundary colored green is $\partial_{L}I$ connecting the minima shown as green points and the part in purple is $\partial_{U}I$ connecting the maxima shown as purple points. The boundary cap is shown in orange which runs parallel to the boundary for a portion of it. Refer to Figure 5 for the zigzag filtration along the boundary cap of the worm. A sequence of graphs with $T=3$ time steps is shown with the corresponding graph filtrations: $\mathcal{F}_{t_{1}},\mathcal{F}_{t_{2}},\mathcal{F}_{t_{3}}$ each with $L=5$ levels. The filtration of the unions are encircled by ovals. Zigzag filtration at the topmost level $\mathcal{Z}_{L}$ is shown in a rectangular box.
@@ 4.2 Computing Zz-Gril
4.3 Time Complexity Analysis
Let $N$ denote the length of the zigzag filtration at the top-most scale level, i.e., ($\mathcal{Z}_{L}$). For the analysis, we consider the zigzag filtration to be starting from an empty complex and ending at an empty complex. Assuming we have temporal data for $T$ time steps and that $L<T$ where $L$ denotes the number of different spatial scales, each iteration in the binary search requires $O(N^{\omega})$ time since each iteration of the binary search involves computing zigzag persistence of a zigzag filtration of length $O(N)$. Here $\omega<2.37286$ is the matrix multiplication exponent. Hence, the binary search in total requires $O(N^{\omega}\log T)$ time per center point. If we sample $s$ center points, then the total time complexity would be $O(sN^{\omega}\log T)$.
@@ 5 Experiments
In this section, we report the results of testing Zz-Gril on various datasets. We begin by giving a detailed description about the experimental setup. Then, we give a brief description of the datasets, followed by the experimental results. We use the benchmark UEA multivariate time-series Bagnall et al. (2018) datasets to test Zz-Gril on multivariate time series data to show that Zz-Gril can be applied to datasets from various domains. Further, we test Zz-Gril on a targeted application of sleep-stage classification by performing experiments on ISRUC-S3 Khalighi et al. (2016) dataset. In all these experiments, we augment the topological information captured by Zz-Gril to one of the specifically tailored machine learning methods on the respective datasets and compare. For each case, we select the machine learning model which has the highest performance to truly test and highlight the value of the topological information added by Zz-Gril to an already high-performing specifically tailored model.
@@ 5.1 Experimental Setup
Given a sample of multivariate time-series data with $m$ time-series, we splice each time-series into a sequence of time-series, each of length $w$. This splicing is done by a moving window of width $w$, where consecutive windows have an overlap of $\lambda$. We calculate the Pearson correlation coefficient Pearson & Lee (1903) between time-series in each window. We construct a graph for each window, where each time-series (of length $w$) is a node, and edges with the Pearson correlation values as weights connect them. Thus, we get a complete graph with $m$ nodes. We select the top $k$ percentile of these edges to build the final graph in each window. This way we obtain a sequence of graphs. The topological information of this sequence of graphs encodes the evolution of correlation between time series. Alternately, we can also track the evolving time-series by converting it into a sequence of point clouds. From the sequence of time-series described above, we consider each time-series of length $w$ as a point in $\mathbb{R}^{w}$. Thus, if we have $m$ time series, we have $m$ points in $\mathbb{R}^{w}$ in each window. Thus, we get a sequence of point clouds in $\mathbb{R}^{w}$. This sequence of point clouds tracks the evolution of each time-series, and a Vietoris-Rips like construction on this point cloud tracks the evolution of the interaction between time-series. Refer to Figure 6 for an illustration.
@@ 5.2 UEA Multivariate Time Series Classification
UEA Multivariate Time Series Classification (MTSC) Bagnall et al. (2018) archive comprises of real-world multivariate time series data and is a widely recognized benchmark in time series analysis. The UEA MTSC collection encompasses a diverse range of application domains such as healthcare (ECG or EEG data), motion recognition (recorded using wearable sensors). See Table 8 in Appendix B.
The datasets are preprocessed and split into training and testing sets. For this set of experiments, we convert the multivariate time series into a sequence of point clouds and compute Zz-Gril. We choose 5 datasets which have at least 7 multivariate time series. This ensures that each point cloud, in the sequence of point clouds, has at least 7 points, giving meaningful topological information. Refer to Figure 7 in Appendix B for a visualization of Zz-Gril on FingerMovements dataset.
In Table 2, we compare the performance of augmenting Zz-Gril to two specifically tailored machine learning models for multivariate time series classification on UEA MTSC datasets, TapNet Zhang et al. (2020) and TodyNet Liu et al. (2024). We can see from the table that Zz-Gril is adding meaningful topological information on most datasets which is seen as an improvement in performance. We picked these two models for augmenting Zz-Gril because of two reasons: (i) availability and ease-of-use of codebase, (ii) these models already have good performance on UEA MTSC Datasets and we wanted to test the additional value of the topological information that Zz-Gril adds. In order to do a comprehensive evaluation of Zz-Gril framework, we compare Zz-Gril in isolation with other methods Li et al. (2021); Tang et al. (2022); Karim et al. (2019); Schäfer & Leser (2017) on UEA MTSC task. We report the results in Table 1. We see that Zz-Gril has comparable performance. Here, we would like to remind the reader that Zz-Gril is a general framework which does not need to be tailored for these datasets in particular. Further, to show that we need both the spatial and the temporal information to be captured together, we compare Zz-Gril with standard zigzag persistence at various spatial scales. We report the results in Table 3. We can see that Zz-Gril performs better than standard zigzag persistence even across different scales.
@@
We conduct two ablation studies. First, we find that converting time series into either, a sequence of point clouds or a sequence of graphs, is viable for extracting topological information as observed in Table 4). Second, to highlight the value of dynamic topological information, we show that Zz-Gril outperforms a “Snapshot PH” baseline, which vectorizes individual non-zigzag persistence modules. As shown in Table 5, Zz-Gril’s superior performance confirms the importance of the features captured by zigzag persistence.
Dataset TodyNet + Zz-Gril point clouds TodyNet + Zz-Gril graphs
FingerMovements 0.660 0.680
NATOPS 0.961 0.945
SelfRegulationSCP2 0.600 0.594
Dataset Snapshot PH Zz-Gril
FingerMovements 0.480 0.590
Heartbeat 0.639 0.721
MotorImagery 0.560 0.580
NATOPS 0.777 0.850
SelfRegulationSCP2 0.466 0.522
Table 5: Comparison of vectorizing each non-zigzag filtration individually (Snapshot PH) versus Zz-Gril.
We report the computation times in Table 6 which indicate that Zz-Gril is, indeed, practical to use.